OpenStudy (anonymous):

Solve: -20 − 3a = -5(a − 2) (a) -15 (b) -19 (c) 15 (d) 19

OpenStudy (anonymous):

-20-3a=-5a+10 as '-' * '-' = '+' soit becom +10 -20-3a+5a-10=0 now simplify it..by adding similar terms

OpenStudy (anonymous):

I got a.. I think I did it wrong

OpenStudy (anonymous):

and is 15 check it wd ur ans.